# Direct CKM determination from W decays at future lepton colliders

### v1

@@ 1 Introduction
A complementary measurement of CKM matrix elements, free from hadronic matrix elements uncertainties, can be performed at future lepton colliders, such as FCC-ee, ILC, and CepC, from the direct on-shell production of $W$ bosons and subsequent decay in two quark jets Charles:2020dfl ; deBlas:2024bmz . The decay width of the process $W^{+}\to u_{i}\bar{d}_{j}$, as well as its conjugate, is directly proportional to the CKM element $|V_{ij}|^{2}$. Colliding electrons and positrons at center of mass energies close to the $WW$ production threshold, these colliders will provide a large sample of events in an exquisitely clean environment. Additionally, state-of-the-art jet flavor taggers, calibrated precisely thanks to the $\approx 10^{12}$ $Z$ bosons produced in the $Z$-pole run, allow to reconstruct the initial quark flavors of an event with high efficiency and relatively low mistag probabilities. For the remainder of this study we focus on the projected performances of jet-flavor taggers at the FCC-ee future collider, obtained with the IDEA detector concept Bedeschi:2022rnj ; Gouskos:taggers ; Selvaggi:taggers , and perform our analysis based on these values. However, we expect similar results to apply also for CepC CEPCPhysicsStudyGroup:2022uwl ; Liang:2023yyi . At the ILC, while a similar number of $W$ bosons can be expected, the lack of a Tera-$Z$ run does not allow to reach the same level of precision in the jet tagging performance ILDConceptGroup:2020sfq . A measurement of this kind for $|V_{cs}|$ was performed at LEP2 DELPHI:1998hlc , reaching a precision of $\sim 16\%$.
@@
$|V_{ij}|$ Current FCC-ee FCC-ee
(PDG) ($\delta_{\epsilon}=1\%$) ($\delta_{\epsilon}=0.1\%$)
$|V_{cs}|$ $0.975\pm 0.006$ $(0.6\%)$ $0.36\%$ $0.05\%$
$|V_{cb}|$ $(40.8\pm 1.4)\times 10^{-3}$ $(3.4\%)$ $0.52\%$ $0.16\%$
Table 1: The firs column shows the current values of $|V_{cs}|$ and $|V_{cb}|$ from PDG ParticleDataGroup:2022pth . The second and third columns show the relative precision projected for FCC-ee by assuming 1% and 0.1% relative systematic uncertainty on the tagger parameters, respectively.
Early studies Charles:2020dfl estimated, from the projected statistics, a relative precision of approximately $0.4\%$ on $|V_{cb}|$, using ILD jet tagging performances as reference. This was later revised to $0.15\%$ using FCC-ee performances Monteil:VcbWW2024 . Our work follows the same direction, improving the previous ones by discussing possible sources of background and considering the effect of systematic uncertainties on tagger performances. Our results for the expected sensitivity from $WW$ decays, for two different values of relative systematic uncertainty $\delta_{\epsilon}=1\%$ and $0.1\%$, are summarized in Table 1 and compared with current values. Such a measurement of CKM matrix elements would be completely independent from those obtained from meson decays. In particular, these will not require any lattice QCD inputs and could provide a final answer to the longstanding tension between inclusive and exclusive determinations of $|V_{cb}|$.
@@ 2 Signal
where $\Gamma_{0}\approx g_{2}^{2}m_{W}/(48\pi)$, $g_{2}$ is the ${\rm SU}(2)_{L}$ coupling constant, $m_{W}=80.4$ GeV is the $W$ boson mass, and the factor of 3 comes from the number of quark colors. The indices run over the kinematically allowed quark flavors: $i=1,2$ for up-quarks and $j=1,2,3$ for down quarks. The same expression is valid for $W^{-}\to\bar{u}_{i}d_{j}$.
@@
${\cal B}_{ij}=\frac{\Gamma_{ij}^{\pm}}{\Gamma_{\rm tot}}=\frac{|V_{ij}|^{2}}{\sum_{lm}|V_{lm}|^{2}}\mathcal{B}_{\rm had}\,.$ (3)
@@
The CKM element $|V_{ij}|$ can be measured by counting the number of decays into the final state $\left(u_{i}d_{j}\right)$, where we are considering both charge combinations. The couple of jets can be produced by either one $W$ boson, while the other produces a different pair of jets, or by both $W$ bosons at the same time. The efficiency of this counting depends on the performance of the jet flavor taggers, which are discussed in Section 4.
QCD corrections can affect the final $4q$ state in several ways. Corrections to the $W\to q\bar{q}^{\prime}$ decay cancel, up to negligible kinematical effects, when we write the branching ratio ${\cal B}_{ij}$ in terms of the measured total hadronic branching ratio as in Eq. (3). On the other hand, QCD interactions between the two pairs of quarks can induce colour reconnection effects, affecting the final hadron distribution Gustafson:1988fs ; Sjostrand:1993hi ; ALEPH:2013dgf . This phenomenon takes place during the fragmentation process and can be modelled by modern showering algorithms Christiansen:2015yca , which is a requirement for a precise measurement of the $W$ boson mass. Since the goal of this work is to provide an estimate of the sensitvity reach in the CKM matrix element extractions, in the following we work at the parton level doing a simple counting analysis and we assume that any uncertainties due to the colour reconnection modelling can be embedded in the systematic uncertainties of the jet flavour tagging efficiencies.
Finally, we note that the semileptonic channel, $e^{+}e^{-}\to WW\to\left(\ell\nu\right)\left(ud\right)$, could also be exploited for the CKM measurement. The main advantage would be the absence of Drell-Yan backgrounds, which we discuss in the next Section, and of colour reconnection effects. On the downside, the smaller branching ratio in leptons leads to a factor of $\sim 1/2$ in the total number of hadronic events, while the reconstruction efficiency is reduced due to the invisible neutrino. Nevertheless, the semileptonic channel can be employed to expand the present study, or to control the tagger efficiencies in the $WW$ runs. We leave these considerations for future works and proceed here to consider the purely hadronic decays.
@@ 3 Background
The main background to the $e^{+}e^{-}\to W^{+}W^{-}\to 4j$ process is the Drell-Yan induced one, $e^{+}e^{-}\to q\bar{q}+2j$, where the two extra jets can be either both gluons or a quark anti-quark pair. To estimate its importance we simulate $10^{6}$ Drell-Yan (DY) induced events at LO with MadGraph5_aMC Alwall:2011uj , restricting for simplicity to parton level. We fix the center of mass energy to $\sqrt{s}=162$ GeV and impose the following cuts: $p_{T}^{j}>5$ GeV, $\eta_{j}<2$, $\Delta R_{jj}>0.1$, yielding the total cross section $\sigma\approx 22.7$pb.
We apply a simple selection cut to the simulated events, by requiring events in which there exist two pairs of jets with invariant masses $(m_{12},m_{34})$ within 5 GeV of $m_{W}$, consistent with the signal being four jets coming from two $W$ decays. We find that $\sim 6\times 10^{3}$ background events pass this cut, with the resulting cross section $\sigma_{4j-m_{W}}\approx 159$ fb. The latter can be decomposed in two categories: final state with two quarks and two gluons, with $\sigma_{2q2g}\approx 158$ fb, and final state with four quarks, with $\sigma_{4q}\approx 0.95$ fb. The vast majority of the brackground events thus falls in the first category, while only $\sim 1\%$ goes in the second. The detailed flavour decomposition of these background events is reported in Table 2. To derive a corresponding number of events, the reported cross section should be multiplied by the integrated luminosity at the $WW$ threshold scan $\mathcal{L}_{WW}\approx 120{\rm ab}^{-1}$ FCC:2018evy . Given the dominance of gluons in the background, we expect an increased rejection rate due to the low mistag probabilities of b-, c-, and s-flavor taggers versus light jets, see Table 4.
    Parameter     Value     Rel. Unc. (%)
$N_{WW}$ $3\times 10^{8}$ $0.2$
${\rm Br}(W\to{\rm had})$ $0.6741$ $0.4-0.01$
${\rm Br}(W\to\ell\nu)$ $0.3278$ $0.9-0.03$
${\cal A}_{W}$ 0.9
Table 3: Nuisance parameters considered in the fit. Relative uncertainties are shown in the last column FCC:2018byv . For $W$ decay branching ratios, both present and FCC-ee projected uncertainties are listed.
@@ 4 Probabilistic model and systematics
 b  s  c  u  d  g
$\epsilon_{\beta}^{b}$ 0.8 0.0001 0.003 0.0005 0.0005 0.007
$\epsilon_{\beta}^{c}$ 0.02 0.008 0.8 0.01 0.01 0.01
$\epsilon_{\beta}^{s}$ 0.01 0.9 0.1 0.3 0.3 0.2
Table 4: Jet-flavor taggers Working Points.
The nuisance parameters $\nu$ include physics and detector inputs, like branching ratios and acceptances, as well as tagger performance parameters. We list the former in Table 3, showing their numerical values and relative uncertainties FCC:2018byv . The latter are shown in Table 4: for the $q$-tagger, we denote as $\epsilon_{\beta}^{q}$ the probability to tag a $\beta$-jet as a $q$-jet, where $\beta=\{b,s,c,u,d,g\}$. Currently, the level of systematic uncertainty on these parameters, $\delta_{\epsilon}\equiv\delta\epsilon/\epsilon$, is of the order of few percent ATLAS:2019bwq ; CMS:2017wtu , while future colliders will be able to reduce it to ${\cal O}(1)\%$. Recent studies on $b$-taggers at FCC-ee Selvaggi:Rb show that a systematic of ${\cal O}(0.1)\%$ is achievable, in particular thanks to the dedicated calibration during the running at the $Z$-pole. In order to show the effect of the systematic uncertainties on $\epsilon_{\beta}^{q}$ in our results, we fix three benchmarks for our numerical analysis: $\delta_{\epsilon}=0,~{}0.1,~{}1$. In each of these cases, we assume for simplicity that $\delta_{\epsilon}$ is the same for all tagger parameters.
Given the discussion in Section 3, we are also assuming that the DY background can be effectively reduced to a negligible fraction, thus not entering in the definition of $N_{B_{bc;1},B_{bc;2}}(\mu,\nu)$. [^2] We checked that including DY events in the case $\delta_{\epsilon}=0$ has a negligible impact in the sensitivity to $|V_{cs}|$ and $|V_{cs}|$. An even smaller effect is expected when systematic uncertainties are included. As a further simplification in our analysis, we exploit the large statistics expected in each tag bin $B_{bc;i}$ to bypass the need of detailed MonteCarlo signal simulations, and instead work in the Asimov approximation Cowan:2010js . In the Asimov approximation, one sets the observed values of $N_{B_{bc;i}}$ to $N^{A}_{B_{bc;i}}=N_{B_{bc;i}}(\mu=\mu_{\rm true},\nu=\nu_{0})$, where $\mu_{\rm true}$ is the input value of $\mu$ (the PDG central value in this work) and $\nu_{0}$ is the nominal value of each nuisance parameter.
@@ 5 Results
Here we apply our method to study the cases of $|V_{cb}|$ and $|V_{cs}|$. The main results are summarized in the last two columns of Table 1 and are obtained by considering one matrix element at a time and only the two more relevant taggers, whose efficiencies are set to the values shown in Table 4. The two columns correspond to two benchmarks for systematic uncertainties, $\delta_{\epsilon}=1\%,~{}0.1\%$. In all cases, FCC-ee is able to improve on the current levels of precision. Here and in the following, we only consider systematic uncertainties for the tagger efficiencies. We have found that the other systematic uncertainties are subleading and do not qualitatively alter our results.
@@
We now move to study the dependence of our results on the specific working point for the tagging efficiencies, for the three benchmarks of the systematic uncertainties. In Fig. 3 (left column) we show isocontours of $\delta|V_{cb}|/|V_{cb}|$ by varying different combinations of tagger parameters. In both plots, we show with black, red and green lines the results with the three different uncertainty assumptions, statistics, 1% and 0.1% systematics respectively, while the red dot indicates the tagger working point values. In the top plot, we scan the $\{\epsilon_{b}^{b},\epsilon_{c}^{c}\}$ plane. The results present only a mild dependence on these parameters: to drop the precision by a factor of 2 for small systematics, the efficiencies of both taggers need to be halved. The same corresponds to a ${\cal O}(10)\%$ drop for the 1% benchmark.
@@
Given the result discussed above and the relatively small statistics of $cb$ final states, the main effect on the precision will come from mistagging a light jet as a $c$- or a $b$-jet. Thus, in the bottom plot we turn our attention to the $\{\epsilon^{b}_{udsg},\epsilon^{c}_{udsg}\}$ plane. It is evident in this scan how a powerful rejection of light jets from the $b$-tagger, together with low systematics, are necessary to obtain the desired sensitivity. While projected performances and uncertainties are promising, it is worth noting that $\epsilon_{udsg}^{b}=0.005$ and $\delta\epsilon/\epsilon=1\%$ will bring the FCC-ee sensitivity on $|V_{cb}|$ down to the same order as the current levels. On the other hand, the requirements on the $c$-tagger performance can remain relatively loose, given the very large number of $c$ quark produced.
@@
Figure 3: Precision contours on when scanning over tagger parameters. The black, red and green lines indicate the results with statistics, 1% and 0.1% systematics uncertainties respectively, while the red dot corresponds to the values of the working point, see Table 4. Left column: precision contours on $|V_{cb}|$ in the $\{\epsilon_{b}^{b},\epsilon_{c}^{c}\}$ (top) and $\{\epsilon_{b}^{udsg},\epsilon_{c}^{udsg}\}$ (bottom) planes. Right column: precision contours on $|V_{cs}|$ in the $\{\epsilon_{c}^{c},\epsilon_{s}^{s}\}$ (top) and $\{\epsilon_{c}^{c},\epsilon_{s}^{udg}\}$ (bottom) planes.
@@
Regarding $|V_{cs}|$, the large effect from systematic uncertainties on $\epsilon$, observed in Fig. 2, is evident in Fig. 3 (right column), where we show isocontours of $\delta|V_{cs}|/|V_{cs}|$ by again varying different combinations of tagger parameters. In the top plot, we show the percentual precision as a function of the two relevant true positive rates, $\{\epsilon_{c}^{c},\epsilon_{s}^{s}\}$. For statistics only, the behaviour is fairly straightforward. Increasing either efficiency increases the precision, and the isocontours are obtained by decreasing $\epsilon_{s}^{s}$ as $\epsilon_{c}^{c}$ increases. The introduction of systematic uncertainties for the tagger efficiencies not only degrades the overall performance but also changes the qualitative features of the scans: the isocontours are almost independent on $\epsilon_{c}^{c}$. This indicates a relatively large lack of sensitivity to changes in $\epsilon_{c}^{c}$, with $\epsilon_{s}^{s}$ driving the precision. A similar behavior is observed in the bottom plot, which shows the scan on the $\{\epsilon_{c}^{c},\epsilon_{udg}^{s}\}$ space.

### v2

@@ 1 Introduction
A complementary measurement of CKM matrix elements, free from hadronic matrix elements uncertainties, can be performed at future lepton colliders, such as FCC-ee, ILC, and CepC, from the direct on-shell production of $W$ bosons and subsequent decay in two quark jets Charles:2020dfl ; deBlas:2024bmz . The decay width of the process $W^{+}\to u_{i}\bar{d}_{j}$, as well as its conjugate, is directly proportional to the CKM element $|V_{ij}|^{2}$. Colliding electrons and positrons at center of mass energies close to the $WW$ production threshold, these colliders will provide a large sample of events in an exquisitely clean environment. Additionally, state-of-the-art jet flavor taggers, calibrated precisely thanks to the $\approx 10^{12}$ $Z$ bosons produced in the $Z$-pole run, allow to reconstruct the initial quark flavors of an event with high efficiency and relatively low mistag probabilities. For the remainder of this study we focus on the projected performances of jet-flavor taggers at the FCC-ee future collider, obtained with the IDEA detector concept Bedeschi:2022rnj ; Gouskos:taggers ; Selvaggi:taggers , and perform our analysis based on these values. However, we expect similar results to apply also for CepC CEPCPhysicsStudyGroup:2022uwl ; Liang:2023yyi . At the ILC, while a similar number of $W$ bosons can be expected, the lack of a Tera-$Z$ run does not allow to reach the same level of precision in the jet tagging performance ILDConceptGroup:2020sfq . A measurement of this kind for $|V_{cs}|$ was performed at LEP2 DELPHI:1998hlc , reaching a precision of $\sim 30\%$. The measurement of the total hadronic branching ratio of the $W$ boson, together with precise SM predictions, can also be used to extract a value for $|V_{cs}|$ with an uncertainty of approximately $1\%$, by using the world averages for $\alpha_{s}$ and the other CKM elements dEnterria:2016rbf ; dEnterria:2020cpv .
@@
$|V_{ij}|$ Current FCC-ee FCC-ee FCC-ee
(PDG) ($\delta_{\epsilon}=1\%$) ($\delta_{\epsilon}=0.1\%$) (Stat. only)
$|V_{cs}|$ $0.975\pm 0.006$ $(0.6\%)$ $0.36\%$ $0.05\%$ $0.008\%$
$|V_{cb}|$ $(40.8\pm 1.4)\times 10^{-3}$ $(3.4\%)$ $0.52\%$ $0.16\%$ $0.14\%$
Table 1: The first column shows the current values of $|V_{cs}|$ and $|V_{cb}|$ from PDG ParticleDataGroup:2022pth . The second, third and fourth columns show the relative precision projected for FCC-ee by assuming 1%, 0.1% and zero relative systematic uncertainty on the tagger parameters, respectively.
Early studies Charles:2020dfl estimated, from the projected statistics, a relative precision of approximately $0.4\%$ on $|V_{cb}|$, using ILD jet tagging performances as reference. This was later revised to $0.15\%$ using FCC-ee performances Monteil:VcbWW2024 . Our work follows the same direction, improving the previous ones by discussing possible sources of background and considering the effect of systematic uncertainties on tagger performances. Our results for the expected sensitivity from $WW$ decays, for two different values of relative systematic uncertainty $\delta_{\epsilon}=1\%$ and $0.1\%$ and for statistical uncertainty only, are summarized in Table 1 and compared with current values. The statistical-only precision for $|V_{cb}|$ is slightly better than the reported value in Ref. Monteil:VcbWW2024 in spite of imperfect tagging due to three times more $WW$ pairs being considered. Such a measurement of CKM matrix elements would be completely independent from those obtained from meson decays. In particular, these will not require any lattice QCD inputs and could provide a final answer to the longstanding tension between inclusive and exclusive determinations of $|V_{cb}|$.
@@ 2 Signal
where at leading order (LO) $\Gamma_{0}\approx g_{2}^{2}m_{W}/(48\pi)$, $g_{2}$ is the ${\rm SU}(2)_{L}$ coupling constant, $m_{W}=80.4$ GeV is the $W$ boson mass, and the factor of 3 comes from the number of quark colors. The indices run over the kinematically allowed quark flavors: $i=1,2$ for up-quarks and $j=1,2,3$ for down quarks. The same expression is valid for $W^{-}\to\bar{u}_{i}d_{j}$.
@@
${\cal B}_{ij}=\frac{\Gamma_{ij}^{\pm}}{\Gamma_{\rm tot}}=\frac{|V_{ij}|^{2}}{\sum_{lm}|V_{lm}|^{2}}\mathcal{B}_{\rm had}\,,$ (3)
where we take the total hadronic branching ratio from the PDG ParticleDataGroup:2022pth combination of LEP and CMS results; note however that, at the time FCC-ee or other future circular lepton colliders will be running, the best measurement will be obtained from that same collider as well.
@@
The CKM element $|V_{ij}|$ can be measured by counting the number of decays into the final state $\left(u_{i}d_{j}\right)$, where we are considering both charge combinations. In order to cancel systematic uncertainties associated with the total rate, the probabilistic model introduced in Section 4 is defined in terms of single-event probabilities per final state and thus constrains $|V_{ij}|$ through the ratio
$\frac{{\cal B}_{ij}}{\mathcal{B}_{\rm had}}=\frac{\Gamma^{\pm}_{ij}}{\sum_{lm}\Gamma^{\pm}_{lm}}=\frac{|V_{ij}|^{2}}{\sum_{lm}|V_{lm}|^{2}}.$ (7)
where the denominator is computed using the PDG ParticleDataGroup:2022pth reported best-fit values for all the CKM matrix elements that we are not leaving free in the fit. In each final state $\left(u_{i}d_{j}\right)$, the couple of jets can be produced by either one $W$ boson, while the other produces a different pair of jets, or by both $W$ bosons at the same time. The efficiency in each final state depends on the performance of the jet flavor taggers, which are discussed in Section 4.
QCD corrections can affect the final $4q$ state in several ways. Higher-order corrections to the $W\to q\bar{q}^{\prime}$ decay cancel, up to negligible kinematical effects, when we write the branching ratio ${\cal B}_{ij}$ in terms of the measured total hadronic branching ratio as in Eq. (3) dEnterria:2016rbf . On the other hand, QCD interactions between the two pairs of quarks can induce colour reconnection effects, affecting the final hadron distribution Gustafson:1988fs ; Sjostrand:1993hi ; ALEPH:2013dgf . This phenomenon takes place during the fragmentation process and can be modelled by modern showering algorithms Christiansen:2015yca , which is a requirement for a precise measurement of the $W$ boson mass. Since the goal of this work is to provide an estimate of the sensitvity reach in the CKM matrix element extractions, in the following we work at the parton level doing a simple counting analysis and we assume that any uncertainties due to the colour reconnection modelling can be embedded in the systematic uncertainties of the jet flavour tagging efficiencies.
Finally, we note that the semileptonic channel, $e^{+}e^{-}\to WW\to\left(\ell\nu\right)\left(u_{i}d_{j}\right)$, could also be exploited for the CKM measurement. The main advantage would be the absence of Drell-Yan backgrounds, which we discuss in the next Section, and of colour reconnection effects. On the downside, the smaller branching ratio in leptons leads to a factor of $\sim 1/2$ in the total number of hadronic events, while the invisible neutrino in the final state might pose other experimental challenges. Nevertheless, the semileptonic channel can be employed to expand the present study, or to control the tagger efficiencies in the $WW$ runs. We leave these considerations for future works and proceed here to consider the purely hadronic decays.
@@ 3 Background
The main background to the $e^{+}e^{-}\to W^{+}W^{-}\to 4j$ process is the Drell-Yan induced one, $e^{+}e^{-}\to q\bar{q}+2j$, where the two extra jets can be either both gluons or a quark anti-quark pair. To estimate its importance we simulate $10^{6}$ Drell-Yan (DY) induced events at LO with MadGraph5_aMC Alwall:2011uj , restricting for simplicity to parton level. We fix the center of mass energy to $\sqrt{s}=162$ GeV and impose the following cuts: $p_{T}^{j}>5$ GeV, $\eta_{j}<2$, $\Delta R_{jj}>0.1$, yielding the total cross section $\sigma\approx 22.7$pb. Through the same process, we find that other backgrounds, as the $e^{+}e^{-}\to Z^{*}Z^{*}/\gamma^{*}\gamma^{*}\to 4j$ chain, are negligible.
We apply a simple selection cut to the simulated events, by requiring events in which there exist two pairs of jets with invariant masses $(m_{12},m_{34})$ within 5 GeV of $m_{W}$, consistent with the signal being four jets coming from two $W$ decays. We find that $\sim 6\times 10^{3}$ background events pass this cut, with the resulting cross section $\sigma_{4j-m_{W}}\approx 159$ fb. The latter can be decomposed in two categories: final state with two quarks and two gluons, with $\sigma_{2q2g}\approx 158$ fb, and final state with four quarks, with $\sigma_{4q}\approx 0.95$ fb. The vast majority of the brackground events thus falls in the first category, while only $\sim 1\%$ goes in the second. The detailed flavour decomposition of these background events is reported in Table 2. To derive a corresponding number of events, the reported cross section should be multiplied by the integrated luminosity at the $WW$ threshold scan $\mathcal{L}_{WW}\approx 12{\rm ab}^{-1}$ FCC:2018evy . Given the dominance of gluons in the background, we expect an increased rejection rate due to the low mistag probabilities of b-, c-, and s-flavor taggers versus gluon jets, see Table 3.
 $b$  $s$  $c$  $u$  $d$  $g$
$\epsilon_{\beta}^{b}$ 0.8 0.0001 0.003 0.0005 0.0005 0.007
$\epsilon_{\beta}^{c}$ 0.02 0.008 0.8 0.01 0.01 0.01
$\epsilon_{\beta}^{s}$ 0.01 0.9 0.1 0.3 0.3 0.2
Table 3: Jet-flavor taggers Working Points, indicated by the probabilities $\epsilon_{\beta}^{q}$ to tag a $\beta$-jet as a $q$-jet, with $q=b,c,s$ and $\beta=b,s,c,u,d,g$ (see text for details). Values are taken from Refs. Bedeschi:2022rnj ; Gouskos:taggers ; Selvaggi:taggers .
@@ 4 Probabilistic model and systematics
    Parameter     Value
$N_{WW}$ $3\times 10^{8}$
${\rm Br}(W\to{\rm had})$ $0.6741$
${\cal A}_{W}$ 0.9
Table 4: Fixed parameters from physics and detector inputs considered in the fit.
The nuisance parameters $\nu$ include physics and detector inputs, like branching ratios and acceptances, as well as tagger performance parameters. We list the former in Table 4, showing their numerical values; their relative uncertainties are small enough FCC:2018byv that their effect on the fit is negligible, thus we consider them as fixed inputs. The latter are shown in Table 3: for the $q$-tagger, we denote as $\epsilon_{\beta}^{q}$ the probability to tag a $\beta$-jet as a $q$-jet, where $\beta=\{b,s,c,u,d,g\}$. Currently, the level of systematic uncertainty on these parameters, $\delta_{\epsilon}\equiv\delta\epsilon/\epsilon$, is of the order of few percent ATLAS:2019bwq ; CMS:2017wtu , while future colliders will be able to reduce it to ${\cal O}(1)\%$. Recent studies on $b$-taggers at FCC-ee Selvaggi:Rb show that a systematic of ${\cal O}(0.1)\%$ is achievable, in particular thanks to the dedicated calibration during the running at the $Z$-pole. In order to show the effect of the systematic uncertainties on $\epsilon_{\beta}^{q}$ in our results, we fix three benchmarks for our numerical analysis: $\delta_{\epsilon}=0,~{}0.1,~{}1$. In each of these cases, we assume for simplicity that $\delta_{\epsilon}$ is the same for all tagger parameters.
Given the discussion in Section 3, we are also assuming that the DY background can be effectively reduced to a negligible fraction, thus not entering in the definition of $N_{B_{bc;1},B_{bc;2}}(\mu,\nu)$. [^2] We checked that including DY events in the case $\delta_{\epsilon}=0$ has a negligible impact in the sensitivity to $|V_{cs}|$ and $|V_{cb}|$. An even smaller effect is expected when systematic uncertainties are included. As a further simplification in our analysis, we exploit the large statistics expected in each tag bin $B_{bc;i}$ to bypass the need of detailed Monte Carlo signal simulations, and instead work in the Asimov approximation Cowan:2010js . In the Asimov approximation, one sets the observed values of $N_{B_{bc;i}}$ to $N^{A}_{B_{bc;i}}=N_{B_{bc;i}}(\mu=\mu_{\rm true},\nu=\nu_{0})$, where $\mu_{\rm true}$ is the input value of $\mu$ (the PDG central value in this work) and $\nu_{0}$ is the nominal value of each nuisance parameter.
@@ 5 Results
Here we apply our method to study the cases of $|V_{cb}|$ and $|V_{cs}|$. The main results are summarized in the last three columns of Table 1 and are obtained by considering one matrix element at a time and only the two more relevant taggers, whose efficiencies are set to the values shown in Table 3. The second and third columns correspond to two benchmarks for systematic uncertainties, $\delta_{\epsilon}=1\%,~{}0.1\%$ and the last column corresponds to the statistical uncertainty only scenario. In all cases, FCC-ee is able to improve on the current levels of precision. Here and in the following, we only consider systematic uncertainties for the tagger efficiencies. We have found that the other systematic uncertainties are subleading and do not qualitatively alter our results.
@@
We now move to study the dependence of our results on the specific working point for the tagging efficiencies, for the three benchmarks of the systematic uncertainties. In Fig. 3 (left column) we show isocontours of $\delta|V_{cb}|/|V_{cb}|$ by varying different combinations of tagger parameters. In both plots, we show with black, red and green lines the results with the three different uncertainty assumptions, statistical uncertainty only, 1% and 0.1% systematics respectively, while the red dot indicates the tagger working point values. In the top plot, we scan the $\{\epsilon_{b}^{b},\epsilon_{c}^{c}\}$ plane. The results present only a mild dependence on these parameters: to drop the precision by a factor of 2 for small systematics, the efficiencies of both taggers need to be halved. The same corresponds to a ${\cal O}(10)\%$ drop for the 1% benchmark.
@@
Given the result discussed above and the relatively small statistics of $cb$ final states, the main effect on the precision will come from mistagging a light jet as a $c$- or a $b$-jet. Thus, in the bottom plot we turn our attention to the $\{\epsilon^{b}_{udsg},\epsilon^{c}_{udsg}\}$ plane. It is evident in this scan how a powerful rejection of light jets from the $b$-tagger, together with low systematics, are necessary to obtain the desired sensitivity. While projected performances and uncertainties are promising, it is worth noting that $\epsilon_{udsg}^{b}=0.005$ and $\delta\epsilon/\epsilon=1\%$ will bring the FCC-ee sensitivity on $|V_{cb}|$ down to the same order as the current levels. On the other hand, the requirements on the $c$-tagger performance can remain relatively loose, given the very large number of $c$ quarks produced.
@@
Figure 3: Precision contours on $|V_{cb}|$ (left) and $|V_{cs}|$ (right) obtained by scanning over tagger parameters. The black, red and green lines indicate the results with statistical uncertainty only, 1% and 0.1% systematics uncertainties respectively, while the red dot corresponds to the values of the working point, see Table 3. Left column: precision contours on $|V_{cb}|$ in the $\{\epsilon_{b}^{b},\epsilon_{c}^{c}\}$ (top) and $\{\epsilon_{b}^{udsg},\epsilon_{c}^{udsg}\}$ (bottom) planes. Right column: precision contours on $|V_{cs}|$ in the $\{\epsilon_{c}^{c},\epsilon_{s}^{s}\}$ (top) and $\{\epsilon_{c}^{c},\epsilon_{s}^{udg}\}$ (bottom) planes.
@@
Regarding $|V_{cs}|$, the large effect from systematic uncertainties on $\epsilon$, observed in Fig. 2, is evident in Fig. 3 (right column), where we show isocontours of $\delta|V_{cs}|/|V_{cs}|$ by again varying different combinations of tagger parameters. In the top plot, we show the percentual precision as a function of the two relevant true positive rates, $\{\epsilon_{c}^{c},\epsilon_{s}^{s}\}$. For statistical uncertainty only, the behaviour is fairly straightforward. Increasing either efficiency increases the precision, and the isocontours are obtained by decreasing $\epsilon_{s}^{s}$ as $\epsilon_{c}^{c}$ increases. The introduction of systematic uncertainties for the tagger efficiencies not only degrades the overall performance but also changes the qualitative features of the scans: the isocontours are almost independent on $\epsilon_{c}^{c}$. This indicates a relatively large lack of sensitivity to changes in $\epsilon_{c}^{c}$, with $\epsilon_{s}^{s}$ driving the precision. A similar behavior is observed in the bottom plot, which shows the scan on the $\{\epsilon_{c}^{c},\epsilon_{udg}^{s}\}$ space.